OpenStudy (anonymous):

cot(theta) is undefined for theta = [check all that apply] A) 0 B) pi/2 C) 3pi/2 D) pi

terenzreignz (terenzreignz):

You'd do well to remember that \[\large \cot\theta = \frac{\cos\theta}{\sin\theta}\] your best bet is finding values of theta for which \(\Large \sin\theta = 0\)

OpenStudy (anonymous):

\[\cot \theta=\frac{ \cos \theta }{ \sin \theta } \] \[It is undefined for which \sin \theta=0,i.e.,\theta=n \pi,where n is an integer.\]